# RPNT: Robust Pre-trained Neural Transformer - A Pathway for Generalized Motor Decoding

### v1

@@ Abstract
Brain decoding aims to interpret and translate neural activity into behaviors. As such, it is imperative that decoding models are able to generalize across variations, such as recordings from different brain sites, distinct sessions, different types of behavior, and a variety of subjects. Current models can only partially address these challenges and warrant the development of pretrained neural transformer models capable to adapt and generalize. In this work, we propose RPNT - Robust Pretrained Neural Transformer, designed to achieve robust generalization through pretraining, which in turn enables effective finetuning given a downstream task. To achieve the proposed architecture of RPNT, we undertook an investigation to determine which building blocks will be suitable for neural spike activity modeling, since components from transformer models developed for other modalities do not transfer directly to neural data. In particular, RPNT unique components include 1) Multidimensional rotary positional embedding (MRoPE) to aggregate experimental metadata such as site coordinates, session name and behavior types; 2) Context-based attention mechanism via convolution kernels operating on global attention to learn local temporal structures for handling non-stationarity of neural population activity; 3) Robust self-supervised learning (SSL) objective with uniform causal masking strategies and contrastive representations. We pretrained two separate versions of RPNT on distinct datasets a) Multi-session, multi-task, and multi-subject microelectrode benchmark; b) Multi-site recordings using high-density Neuropixel 1.0 probes. The datasets include recordings from the dorsal premotor cortex (PMd) and from the primary motor cortex (M1) regions of nonhuman primates (NHPs) as they performed reaching tasks. After pretraining, we evaluated the generalization of RPNT in cross-session, cross-type, cross-subject, and cross-site downstream behavior decoding tasks. Our results show that RPNT consistently achieves and surpasses the decoding performance of existing decoding models in all tasks. Our ablation and sweeping analysis demonstrate the necessity and robustness of the proposed novel components.
1 Introduction
The transformer architecture (Vaswani et al., 2017), along with pretraining, has greatly shifted data modeling paradigms in the field of natural language processing and subsequently in a variety of fields. Indeed, transformer models such as BERT (Devlin et al., 2019), GPT (Radford et al., 2019), and Vision Transformers (Dosovitskiy et al., 2020) demonstrated that large-scale pretraining on diverse data followed by downstream task-specific finetuning yields more enhanced performance on a variety of related tasks compared to task-specific models. However, for neural data, which contains session-based nonstationary spatiotemporal structure, neural models have yet to utilize such strategies in the full extent. Typical variations in neural data are due to recordings from different brain sites, distinct sessions, different types of behavior, and a variety of subjects.
Efforts have been made recently. Methods such as Latent Factor Analysis via Dynamical Systems (LFADS) (Pandarinath et al., 2018) pioneered multi-session pretraining. Later, transformer-based models such as self-supervised Neural Data Transformer and its variants (NDT (Ye and Pandarinath, 2021), NDT2 (Ye et al., 2023), NDT3 (Ye et al., 2025)) and supervised pretraining models such as POYO (Azabou et al., 2023), PoYO++ (Azabou et al., 2024), PoSSM (Ryoo et al., 2025) showed the benefits of pretraining and finetuning in the neural decoding workflow.
While these efforts have adapted pretraining strategies to neural datasets with larger samples of neural activity (i.e., neural population and neural dynamics models) and more diverse samples (distinct recording sessions and subjects), robust generalization via explainable mechanisms for possible variations in neural data remains a challenge. In particular, several key aspects of variation in neural data have yet to be resolved. First, model generalization capabilities remain limited when applied to unseen brain sites or different recording configurations (Jude et al., 2022; Karpowicz et al., 2024; Le et al., 2025). Second, neural signals drift over time (Chestek et al., 2011). Thus, building powerful mechanisms that can handle neural non-stationarity remains unclear. Third, existing approaches typically emphasize training objectives that ultimately improve performance in decoding tasks by denoising neural signals, but may be unable to learn representations that are informative of the underlying causal dynamics that give rise to neural activity (Lu et al., 2025).
Figure 1: Overall illustration of the pretraining and finetuning workflow for generalized motor decoding. (A) Experimental setup for data collection while NHPs performed reaching tasks. (B) Preparation of pretaining data. (C) and (D) overall schemes for SSL and SFT, respectively. (E) Illustration of model adaptation during pretraining and finetuning.
To address the above limitations, we introduce RPNT - a neural transformer model that incorporates robust pre-training enhancements. RPNT contains novel components that are a result of an investigation aimed at designing a transformer model specifically for neural activity and its variations. In particular, the key novel components of RPNT are: 1) Multidimensional rotary positional embedding (MRoPE) to aggregate experimental metadata such as site coordinates, session name and behavior types; 2) Context-based attention mechanism via convolution kernels operating on the global attention to learn local temporal structures for handling the non-stationarity of neural activity; 3) Robust self-supervised learning (SSL) objective with uniform causal masking strategies and contrastive representations.
For validation of RPNT, we implemented a standard pretraining and finetuning workflow (see Figure 1). With these workflows, we pretrained two separate versions of RPNT on two datasets which include different modalities of the recorded neural activity: a) Multi-session, multi-task, and multi-subject microelectrode benchmark data (Perich et al., 2018); b) Multi-site recordings using high-density Neuropixel 1.0 probes (Canfield et al., 2025). Both datasets include recordings from the (PMd) and the (M1) regions of nonhuman primates (NHPs) as they performed center-out or random target reaching tasks. After pretraining, we evaluated the generalization of RPNT in cross-session, cross-type, cross-subject, and cross-site downstream behavior decoding tasks. Our results showed that RPNT consistently achieved and surpassed the decoding performance of existing decoding models in all tasks. Our ablation and sweeping analysis demonstrated the necessity and robustness of the proposed novel components. Furthermore, the model’s learned attention map provided data-driven insights into the underlying structure of neural encoding of movement variables. Our approach has implications for future large-scale, robust neural transformer pretraining for neural decoding applications such as brain-computer interfaces (BCIs) (Fang and Yang, 2022) and data-driven discovery for neuroscience (Canfield et al., 2025). To this end, our major contributions are listed:
  • Model architecture: We propose a multidimensional rotary positional embedding (MRoPE) to aggregate experimental metadata for decoding generalization. We introduce a context-based local attention mechanism to handle the neural non-stationarity.
  • Pretraining strategy: We pretrain RPNT using a robust SSL objective via uniform causal masking strategies and contrastive representations.
  • Generalized Decoding: We demonstrate superior decoding performance for 1) cross-session, cross-behavior types, and cross-subject scenarios in the microelecotrode recording benchmark; 2) cross-site scenario in a new Neuropixels dataset.
2 Related Work
Neural Population Modeling Classical approaches to neural population analysis proposed to leverage dimensionality reduction to extract low-dimensional structure from high-dimensional recordings. Methods such as PCA (Cunningham and Yu, 2014) and Factor Analysis (Santhanam et al., 2009) were applied to neural activity and showed identification of dominant modes of variance. Since such methods rely on the assumption of static and linear relationships, extensions have been developed, such as a probabilistic geometric PCA with application to neural data (Hsieh and Shanechi, 2025) and Canonical Correlation Analysis to cross-region interactions (Semedo et al., 2019). While extending classical dimension reduction approaches, such methodologies remain limited to pairwise correlations. As an alternative, dynamical systems approaches such as GPFA (Yu et al., 2008) and LFADS (Pandarinath et al., 2018) have been introduced to model the temporal evolution of latent states. While such models are applicable to ubiquitous interpretation of neural population data, their main applications are focused on offline neural denoising analysis rather than neural decoding.
@@ 2 Related Work
Motor Decoding Motor decoding methods aim to reconstruct intended movements or kinematic variables from neural population activity. Classical approaches, such as Wiener (Van Drongelen, 2018) and Kalman filters (Wu et al., 2002; Orsborn et al., 2014), provide real-time decoding with computational efficiency, however typically assume linear dynamics and Gaussian noise. Bayesian decoders (Wu et al., 2006; Shanechi et al., 2016) incorporate prior information about movement statistics, improving robustness but remaining limited by parametric assumptions. Deep learning approaches have been developed as well, and these progressed from MLPs (Glaser et al., 2020) to RNN architectures (Mante et al., 2013; Sani et al., 2024) aimed to capture temporal dependencies in neural dynamics and thus perform effective decoding. Recent works explore Transformer-based decoders (Ye and Pandarinath, 2021; Le and Shlizerman, 2022) that leverage self-attention to model long-range dependencies across neural populations. However, these methods predominantly train from scratch on limited task-specific data, leading to poor generalization across sessions, subjects, and experimental conditions.
@@
Pretraining of Neural Models Efforts have been made to use large-scale pretraining for neural data analysis with various model architectures. Before the era of the transformer, RNN-based models were explored for multi-session pretraining, such as MRNN (Sussillo et al., 2016) and LFADS (Pandarinath et al., 2018). Then, transformer architecture-based models, i.e., Neural Data Transformer (NDT) (Ye and Pandarinath, 2021) and Pre-training On manY neurOns (PoYo) (Azabou et al., 2023) work and subsequent works represent additional developments toward pretraining of neural models. In particular, NDT (Ye and Pandarinath, 2021) introduced a transformer architecture for spike train modeling, while NDT2 (Ye et al., 2023) demonstrated multi-context pretraining across sessions, subjects, and tasks using a spatiotemporal transformer. NDT3 (Ye et al., 2025) scaled it to hundreds of datasets and showed emergent zero-shot/few-shot capabilities. Concurrently, supervised pretraining models such as POYO (Azabou et al., 2023), PoYO++ (Azabou et al., 2024), PoSSM (Ryoo et al., 2025) took a complementary approach of leveraging learnable neural token representation to realize scalable pretraining across multiple subjects. Additional approaches have been proposed as well. Models such as Population Transformer (PopT) used standard transformer encoder architecture but with novel ensemble-wise and channel-wise discrimination tasks to train their model (Chau et al., 2025). Further, BrainBERT (Wang et al., 2023) proposed to adapt language model architectures directly, treating neural signals as text sequences. Although effective, these transformer models lack components for dealing with the non-stationarity and recording configurations in session-based neural spike activity.
Attention Mechanisms and Positional Embedding for Non-stationary Data: The originally developed self-attention mechanism in (Vaswani et al., 2017) assumes stationary input distributions, which may not be suitable for neural recordings where signal characteristics drift over time (Chestek et al., 2011). Adaptive attention mechanisms have been proposed for domains with nonstationary data, such as video (Chen et al., 2021). These mechanisms are based on learnable kernels to modulate the attention based on local context. Further, in (Zhu et al., 2021) a sliding window attention was introduced to capture local structure, and in (Yang et al., 2016) a hierarchical attention across multiple timescales was employed. Non-stationary Transformers (Liu et al., 2022) introduce series stationary and nonstationary attention to handle distribution shifts in time-series forecasting. FEDformer (Zhou et al., 2022) employs frequency-enhanced decomposed architectures to capture both seasonal and trend patterns in nonstationary signals. Crossformer (Zhang and Yan, 2023) utilizes dimension-segment-wise attention to capture cross-time and cross-dimension dependencies.
Beyond the attention mechanism, positional encoding also presents challenges due to the unique nature of neural data. Particularly, positional encoding is expected to take into consideration the joint recording configurations and time. While standard sinusoidal encodings (Vaswani et al., 2017) and learned positional embeddings (Devlin et al., 2019) are generally applicable, recent positional encodings could offer advantages for modeling relative position exclusive to one of the dimensions, space or time. In particular, RoPE (Su et al., 2024) allows modeling of relative positions exclusively in the temporal dimension. Recent extensions of RoPE to multiple dimensions have shown promise in handling complex positional relationships beyond simple sequential ordering. The Qwen2 (Wang et al., 2024b) demonstrates effective multi-dimensional RoPE for handling both sequence position and other modality dimensions. Here, we embrace RoPE for neural activity (Azabou et al., 2023) and use the extended MRoPE (Wang et al., 2024b) for capturing recording configurations.
3 Methods
Figure 2: A schematic of components in RPNT. Components in black indicate standard transformer signal flow (i.e, no masking and standard attention mechanism). Our novel proposed components include MRoPE (green), context-based attention (cyan), and uniform random masking strategy (pink). MRoPE incorporates experimental metadata into the transformer positional embeddings. The context-based attention captures local temporal structures. The input spike activity is randomly masked according to a uniform distribution for causal masked neural modeling. The masked input spike activity is projected into neural embeddings, forming Q, K, and V representations. MRoPe is applied to get the rotary queries and keys. Attention scores are modulated by context-based convolutional kernels and combined with causal masking before aggregating with values.
3.1 MRoPE: Multi-dimensional Rotary Positional Embedding
Since neural recordings vary, it is imperative that each neural recording configuration should be explicitly modeled within the transformer positional embedding. Such embedding is beyond the temporal positional embedding. For example, in multiple brain sites recordings, an effective positional embedding would be one that can model both spatial site locations and temporal relationships for the generalization to unseen site configurations. A plausible candidate for such embedding that we found is Rotary Positional Embedding (RoPE) (Su et al., 2024), which we extend here to multiple dimensions, i.e., MRoPE, where $M$ stands for the configuration dimensions, to aggregate experimental metadata. In the Neuropixel dataset, $M=3D$, and it includes two-dimensional spatial coordinates (denoted by $(x_{s},y_{s})$) and time $t$. To construct the rotational matrix, we partition the model dimension ($d_{\text{model}}$) into three groups independently ($d_{\text{coord}}=\frac{d_{\text{model}}}{3}$) to represent dimensions each for $x$-coordinate, $y$-coordinate, and temporal position. For each group, we define rotation frequencies:
where $i\in[0,d_{\text{coord}}/2)$. Spatial dimensions use lower frequency ($f=5000$) for coherence across nearby brain sites, while temporal dimensions use standard RoPE frequency ($f=10000$) for fine-grained temporal resolution.
Rotation Operation: In contrast to traditional positional encodings added to embeddings, MRoPE works similarly to RoPE, by applying rotations directly to query and key vectors during attention computation. For 3D position $(x_{s},y_{s},t)$, we construct a block-diagonal rotation matrix $\mathbf{R}_{\text{3D}}(x_{s},y_{s},t)$(see  equation 2) that independently rotates dimension pairs. The attention score between location $i$ and $j$ follows $\mathbf{q}_{i}^{T}\mathbf{k}_{j}=\mathbf{q}^{T}\mathbf{R}_{\text{3D}}^{T}(x_{i},y_{i},t_{i})\mathbf{R}_{\text{3D}}(x_{j},y_{j},t_{j})\mathbf{k}=f(x_{j}-x_{i},y_{j}-y_{i},t_{j}-t_{i})$, which preserves the key relative position property. Such operations enable zero-shot generalization to arbitrary brain site recording configurations while maintaining rotational invariance (also see the appendixF).
MRoPE is designed for arbitrary cases where the recording configurations encounter different subjects, recording times, locations, and behavior types (see Figure 2). Specifically, we can partition the model dimension $d_{\text{model}}$ into $M$ groups depending on the number of recording configurations, i.e., $\theta^{(1)}_{i}=\frac{1}{f_{1}^{2i/d_{1}}},\quad\cdots\quad,\theta^{(\mathcal{M})}_{i}=\frac{1}{f_{M}^{2i/d_{M}}}$, where $d_{1},\cdots,d_{M}$ are the dimensions for each group that in total contribute to $d_{\text{model}}$. The $f_{1},\cdots,f_{M}$ are the corresponding frequencies. Indeed, for the public benchmark, we set MRoPE with $M=4D$ to represent behavior types, subjects, recording times, and temporal position. We used an equal dimension partition ($d_{m}=\frac{d_{model}}{4}$) and set frequencies with 10, 100, 1000, and 10000, respectively (see the appendixF).
3.2 Context-Based Attention Mechanism
To address non-stationary neural activity (Chestek et al., 2011), we introduce a context-based attention mechanism via learnable convolution kernels (Gulati et al., 2020) operating on global attention to learn the local temporal structure. Such attention is achieved through the incorporation of history and pooling into attention. We also illustrate the attention steps (cyan color) in Figure 2.
Context Generation: Given a batch neural input $\mathbf{x}\in\mathbb{R}^{B\times T\times D}$ and historical data $\mathbf{H}_{\text{hist}}\in\mathbb{R}^{B\times T_{\text{hist}}\times D}$ containing only past timesteps, we generate a context vector through attention pooling:
We note that $T_{\text{hist}}\subseteq T$. In practice, one can simple use $T_{\text{hist}}=T$ for considering up-to-date historical data as the context history for the generation of convolution kernels. Another computationally efficient choice for $T_{\text{hist}}$ is to use receding windows, e.g., we will use the most recent K time bines. In this work, we use the first implementation for simplicity.
Dynamic Kernel Generation: We parameterize the context vector to generate 2D convolution kernels for each attention head:
where $h\in[1,H]$ indexes attention heads and $(K_{1},K_{2})$ are hyperparameters for convolution kernel dimensions.
Kernel-Enhanced Attention: As the last step, we apply a 2D convolution kernel after calculating the standard attention matrix for better capturing the local attention relationship:
where $\mathbf{M}_{\text{causal}}$ is the causal masking ($\mathbf{M}_{\text{causal}}[i,j]=\mathbb{1}[j\leq i]$). This design enables each head to learn specialized temporal dependencies for handling non-stationarity (see Figure 2).
3.3 Neural Transformer Architecture
With the above two components, we compose RPNT’s neural transformer encoder, where we decouple temporal and cross-site processing (for the neuropixel dataset). First, we build a temporal encoder. Given neural spike activity $\mathbf{X}\in\mathbb{R}^{B\times S\times T\times N}$ (batch, sites, time, neurons) and corresponding experimental metadata, we first embed each recording site $\mathbf{H}_{s}^{(0)}=\text{Linear}(\mathbf{X}_{:,s,:,:})$. Then, we build a temporal encoder that consists of $L$ transformer layers with MRoPE and context-based attention
Most of the neural representation learning and neural population modeling can be done via the above powerful temporal encoder by treating the number of recording sites $S=1$, i.e., repeated neural measurements across sessions. Further, if it is desired to model the cross-site neural dynamics beyond the temporal encoding, such as sampled neuropixel recordings for cross-site activity, we propose to concatenate a spatial encoder right after the temporal encoder. Spatial encoder will model cross-site interactions through timestep-wise processing, naturally preserving causality. At each timestep $t$, we follow the standard multihead attention (MHA)
$\mathbf{Z}_{:,:,t,:}^{(l+1)}=\mathbf{Z}_{:,:,t}^{(l)}+\text{MultiheadAttn}(\text{LN}(\mathbf{Z}_{:,:,t,:}^{(l)})),$ (7)
where $\mathbf{Z}^{(0)}=\mathbf{H}^{(L)}$ from the last layer of temporal encoder. The MHA weights can provide interpretable cross-site functional connectivity. In our experiments, we use the microelectrode benchmark (Perich et al., 2018) mainly for the purpose of investigating and validating the power of the temporal encoder and the neuropixel dataset to further show the attention map visualization for data-driven discovery by the spatial encoder.
3.4 Causal Masked Neural Modeling
With the neural transformer encoder setup, we introduce a pretraining framework tailored for neural spike data that robustly performs self-supervised learning (SSL) using masked causal Poisson reconstruction loss with uniform random masking and cross-site contrastive loss.
Uniform Random Masking Strategy. Unlike fixed masking strategies, which required careful tuning in prior work (Srivastava et al., 2014; Ye et al., 2025; Zhang et al., 2024; He et al., 2022), we apply masking both in space (neurons) and time according to the uniform distribution, i.e., $p_{m}\sim\mathcal{U}(0,1)$, at each batch. This choice is motivated by works of  Azabou et al. (2023) and  (Le et al., 2025), where in Azabou et al. (2023) it was shown that dynamic dropout augmented the decoding performance, and in a recent work of (Le et al., 2025) it was shown that uniform neuron sampling strategy can improve the supervised motor decoding performance. These findings led us to further extend masking to both time and neuron during SSL pretraining. The stochastic approach that we propose eliminates the masking rate hyperparameter while ensuring the model encounters diverse reconstruction difficulties during the pretraining phase, i.e., $\mathbf{M}_{:,t,n}\sim\text{Bernoulli}(1-p_{m})$.
Self-Supervised Learning Objective. Distinct from existing denoising-based transformer work, our approach maintains causality, i.e., reconstruction at time $t$ depends only on unmasked inputs from $t^{\prime}\leq t$. This transforms the standard MAE objective He et al. (2022) into a predictive task aligned with the autoregressive nature of neural mechanisms (see Appendix for the comparison). For masked position $(t,n)$, RPNT reconstructs it with respect to
where $\hat{\lambda}_{:,t,n}=\text{Decoder}(\mathbf{Z}_{:,t,d})$ and $\epsilon=10^{-8}$ provides numerical stability (see decoder design in appendix G). Further, an auxiliary objective encourages learning site-invariant representations (see the appendix H.2 for the formulation and ablation study). The combined objective $\mathcal{L}=\mathcal{L}_{\text{recon}}+\mu\mathcal{L}_{\text{contrast}}$ balances reconstruction objective with robust representation learning.
3.5 Downstream Evaluation Framework
$\hat{\mathbf{y}}_{t}=\text{MLP}_{\text{task}}(\mathbf{H}_{:,t}),$ (10)
where $\mathbf{H}_{:,t}$ is the temporal encoder outputs across at time $t$. We use mean-squared error (MSE) for SFT on velocity regression tasks and evaluate using $R^{2}$ metric.
Further, we visualize functional connectivity (FC) based on the spatial attention map across time. The spatial encoder’s cross-site attention weights directly encode functional relationships. We extract the connectivity matrix $\mathbf{C}\in\mathbb{R}^{T\times S\times S}$ by averaging attention weights across layers, i.e., $\mathbf{C}_{i\rightarrow j}(\text{attention})=\frac{1}{L}\sum_{l=1}^{L}\text{Attention}_{i\rightarrow j}^{(l)}$. The visualization results were shown in the appendix J.
4 Experiments
4.1 Datasets and experimental setup
Datasets We evaluated the downstream generalization of the RPNT model under three out of four scenarios of variation (cross-session, cross-task, and cross-subject scenarios) on the public microelectrode benchmark (Perich et al., 2018). For the fourth scenario of cross-site variation, we performed experiments with a Neuropixel dataset. The  (Perich et al., 2018) dataset contains electrophysiology and behavioral data from $4$ rhesus macaques (subjects C, J, M, and T) performing either a Center-Out reaching task (denoted as CO) or a continuous random target acquisition task (denoted as RT). Neural activities were recorded from Utah arrays in the M1 or PMd region. This dataset contained 111 sessions, spanning 43 recording hours. Both raw data and preprocessed data can be accessed from the Python package brainsets provided by (Azabou et al., 2023). In the Neuropixel dataset, a male rhesus macaque (named subject B) was implanted with a multi-modal chamber over the frontal motor cortex. The subject was trained to perform the CO reaching tasks. At each experimental session, a single neuropixel probe was inserted in the chamber to acquire specific recording-site neural activity. The dataset is comprised of 17 recordings at different sites distributed across the PMd and M1 regions across experiment sessions spanning approximately one year (see appendix N).

### v2

@@ Abstract
Brain motor decoding aims to interpret and translate neural activity into behaviors. Decoding models that are able to generalize across variations, such as recordings from different brain sites, experimental sessions, behavior types, and subjects, will be critical for real-world applications. Current decoding models only partially address these challenges. In this work, we develop a pretrained neural transformer model, RPNT - Robust Pretrained Neural Transformer, designed to achieve robust generalization through pretraining, which in turn enables effective finetuning for downstream motor decoding tasks. We achieved the proposed RPNT architecture by systematically investigating which transformer building blocks could be suitable for neural spike activity modeling, since components from models developed for other modalities, such as text and images, do not transfer directly to neural data. The final resulting RPNT architecture incorporates three unique enabling components: 1) Multidimensional rotary positional embedding (MRoPE) to aggregate experimental metadata such as site coordinates, session ids and behavior types; 2) Context-based attention mechanism via convolution kernels operating on global attention to learn local temporal structures for handling non-stationarity of neural population activity; 3) Robust self-supervised learning (SSL) objective with stochastic causal masking strategies and contrastive representations. We pretrained two versions of RPNT on distinct datasets that present significant generalization challenges: a) Multi-session, multi-task, and multi-subject microelectrode benchmark; b) Multi-site recordings using high-density Neuropixel 1.0 probes from many cortical locations. Both datasets include recordings from the dorsal premotor cortex (PMd) and from the primary motor cortex (M1) regions of nonhuman primates (NHPs) performing reaching tasks. After pretraining, we evaluated RPNT generalization in terms of cross-session, cross-type, cross-subject, and cross-site downstream behavior decoding tasks. Our results show that RPNT consistently outperforms the decoding performance of existing decoding models on these tasks. Our ablation and hyperparameter sweep analyses demonstrate the necessity and robustness of the proposed novel components. Our work has implications for future robust brain decoder designs in brain-computer interfaces (BCIs).
Introduction
The transformer architecture [55], together with pretraining, has notably shifted data modeling paradigms, beginning in Natural Language Processing (NLP) and subsequently across many fields. Scalable transformer models such as BERT [15], GPT [41], and Vision Transformers [16] demonstrated that a large scale pretraining on diverse data, followed by task specific finetuning, yields superior performance across related tasks compared with models trained for individual tasks. However, current neural modeling frameworks have not yet fully capitalized on transformer pretraining, in part because neural data exhibit session-dependent and nonstationary spatiotemporal structure. As an alternative, classical approaches leveraged dimensionality reduction to extract low-dimensional structure from high-dimensional session recordings. Methods such as Principal Component Analysis (PCA) [12] and Factor Analysis [46] were applied to neural activity, which captured and identified dominant modes of variance. Such methods assume simple, static, and linear relationships. Thus, extensions have been developed, such as the recent probabilistic geometric PCA method for neural data application [27] and Canonical Correlation Analysis (CCA) for modeling pairwise and cross-region interactions [47]. While these methods extend classical dimension reduction approaches, they remain limited in capturing sequential neural properties. As an alternative, dynamical systems approaches such as Gaussian-Process Factor Analysis (GPFA) [66] model the stochastic temporal evolution of neural latent states. These models have generally been applied to facilitate the interpretation of neural population data in offline analysis applications.
@@ Introduction
In parallel, motor decoding models aim to reconstruct intended movement variables (e.g., kinematics) from neural population activity. Well-established approaches, such as Wiener [54] and Kalman filters [60, 36], provide computationally efficient real time decoding, but typically assume linear stochastic dynamics. Bayesian decoders [61, 48] incorporate prior information about movement statistics, improving robustness but remaining constrained by parametric assumptions. Deep learning approaches have also been developed, progressing from MLPs [22] to RNN architectures [34, 45] that capture temporal dependencies in neural dynamics for effective decoding. More recent work explored Transformer-based decoders [64, 31] that leverage self-attention to model long range dependencies across neural populations. However, these methods are typically trained from scratch on limited task specific data, resulting in poor generalization across sessions, subjects, and experimental conditions. To address this limitation, recent efforts explored large scale pretraining for neural data analysis using a variety of model architectures. Prior to transformer-based models, RNN-based approaches were pioneered and explored for multi-session pretraining, such as Multiplicative Recurrent Neural Network (MRNN) [52] and Latent Factor Analysis via Dynamical Systems (LFADS) [37].
@@
More recent advances in transformer architecture models represent further progress towards pretrained neural models. In particular, Neural Data Transformer (NDT) [64] introduced a transformer architecture for spike train modeling, while NDT2 [63] extended to multi context pretraining across sessions, subjects, and tasks using a spatiotemporal transformer. NDT3 [65] scaled it to hundreds of datasets and showed emergent zero-shot/few-shot capabilities. Concurrently, supervised pretraining models such as Pre-training On manY neurOns (PoYo) [1, 2] and PoSSM [43] adopted a complementary approach that leverages learnable neural token representations to enable scalable pretraining across multiple subjects. Additional approaches, such as Population Transformer (PopT) used a standard transformer encoder architecture but with novel ensemble-wise and channel-wise discrimination tasks for training [6]. BrainBERT [56] further adapts language model architectures by treating neural signals as text sequences. While effective, these transformer models lack explicit mechanisms to address non-stationarity and recording configuration variability in session based neural spike activity.
The aforementioned efforts show the promise of adapting pretraining strategies to neural datasets with larger samples of neural activity (i.e., neural population and neural dynamics models) and more diverse samples (distinct recording sessions and subjects). Yet, robust generalization, via explainable mechanisms for variations in neural data, continues to present challenges, with several key aspects of variation in neural data remaining insufficiently addressed. First, model generalization remains limited when applied to unseen brain regions or different recording configurations [28, 29, 30, 4]. Second, neural signals exhibit drift over time [9]. Decoding efforts will benefit from improved mechanisms to robustly handle neural nonstationarity. Third, existing approaches often emphasize training objectives that improve decoding performance through denoising neural signals, but may fail to learn representations that reflect the underlying causal dynamics that underly neural activity patterns [33]. Decoders that aim to learn causal brain-behavior relationships will likely be beneficial for data-driven discovery and real-time applications.
We introduce RPNT - a neural transformer model that leverages robust pretraining enhancements to address the limitations of current approaches. RPNT contains novel components that emerged from our investigation of transformer architectures tailored for neural activity and its variations. The key novel components of RPNT are: 1) Multidimensional rotary positional embedding (MRoPE) that aggregates experimental metadata such as site coordinates, session ids, and behavior types; 2) A context-based attention mechanism that applies convolution kernels to global attention to learn local temporal structure and handle the non-stationarity of neural activity; 3) A robust self supervised learning (SSL) objective with stochastic causal masking strategies and contrastive representation learning. For experimental validation, we followed and implemented a standard pretraining and finetuning workflow (see Figure 1). Using this workflow, we pretrained two versions of RPNT on two datasets containing different modalities of recorded neural activity: a) A multi-session, multi-task, and multi-subject microelectrode benchmark data [39]; b) A multi-site recordings using high-density Neuropixel 1.0 probes sampling neural activity across many cortical locations [4]. Both datasets include recordings from PMd and M1 regions of nonhuman primates (NHPs) as they performed center-out or random target reaching tasks. After pretraining, we evaluated the generalization of RPNT on cross-session, cross-type, cross-subject, and cross-site behavioral decoding tasks. RPNT consistently matched or exceeded the decoding performance of existing decoding models across all tasks. Ablation and hyperparameter analyses demonstrated the necessity and robustness of the proposed novel components. Furthermore, we observed that the learned spatial attention maps provide data-driven insights into the structure of neural encoding of movement variables. Overall, our investigations and the novel components of RPNT support future large-scale neural transformer pretraining for decoding applications such as brain computer interfaces (BCIs) [42] and data-driven discovery for neuroscience [4]. To this end, our main contributions are:
  • Model architecture: We propose a Multidimensional Rotary Positional Embedding (MRoPE) to aggregate experimental metadata for decoding generalization and introduce a context-based attention mechanism to handle the non-stationarity.
  • Pretraining strategy: We pretrain RPNT using a robust self-supervised learning objective via stochastic causal masking strategies and contrastive representations.
  • Generalized Decoding: We demonstrate superior decoding performance for 1) cross-session, cross-behavior types, and cross-subject scenarios in the microelecotrode recording benchmark; 2) cross-site scenarios in a new Neuropixels dataset.
Figure 1: Overview of the pretraining and finetuning workflow for generalized motor decoding. (A) Experiments in which NHPs performed reaching tasks. (B) Preparation of training, validation, and test splits for the pre-training datasets. session. (C, D) Schematics for self-supervised learning (SSL) and downstream supervised finetuning (SFT), respectively. (E) Illustration of model parameter adaptation during the pretraining and finetuning phases.
Results
Datasets and data collection
Long-term recordings of motor and premotor cortical spiking activity during reaching in monkeys (LTRCH) benchmark dataset We evaluated the downstream generalization capabilities of the RPNT in the presence of three forms of neural variability captured by an existing benchmark dataset: "Long-term recordings of motor and premotor cortical spiking activity during reaching in monkeys" [40] (LTRCH). The dataset allows evaluation of cross-session, cross-task, and cross-subject generalization, as described in prior work [39]. The LTRCH dataset contains electrophysiology and behavioral data from $4$ rhesus macaques (subjects C, J, M, and T) performing a center-out reaching task (denoted as CO) or a continuous random target acquisition task (denoted as RT). Neural activities were recorded from chronically-implanted electrode arrays in the primary motor cortex (M1) or dorsal premotor cortex (PMd). LTRCH contains 111 sessions, spanning 43 recording hours. Both raw data and preprocessed data can be accessed from the Python package brainsets [1].
Neuropixel dataset (NPCS) We evaluated whether RPNT can also generalize across measurement location within the brain (cross-site) using a newly-generated dataset: "NeuroPixels across Cortical Sites" (NPCS) [4]. The dataset allows evaluation of cross-site generalization within a single behavioral task and subject. The NPCS dataset includes electrophysiology and behavioral data from one male rhesus macaque (subject B) performing a center-out reaching task (denoted as CO). In each experimental session, a single neuropixel probe was inserted into the cortex via an implanted recording chamber to measure neural activity from 384 electrodes simultaneously. The probe insertion location was varied across sessions to span over 1 cm of M1 and PMd (see Neuropixel probe setup in Figure 2 and Appendix Table 2). The dataset is comprised of 17 recordings at different recording sites collected during experiment sessions spanning approximately one year (see Appendix 2). The NPCS dataset was collected at the University of Washington and the Washington National Primate Research Center. All procedures were conducted in compliance with the NIH Guide for the Care and Use of Laboratory Animals and were approved by the Institutional Animal Care and Use Committee at the University of Washington. Details of the Neuropixel recording setup and data collections are further described in our prior work [4].
RPNT pretraining and evaluation